The square and the equilateral triangle at the right have the same perimeter.Find the length of the sides of the triangle. One side of the triangle has 3x and the square has a side that is x+5. please help me! I have no clue on how to solve this!
Answer by Earlsdon(6294)

The perimeter of a square is 4*S where S = x+5.
The perimeter of an equilateral triangle is 3*s where s = 3x.
The perimeters are equal, so...
4(x+5) = 3(3x) Simplify.
4x+20 = 9x Subtract 4x from both sides.
20 = 5x Divide both sides by 5.
x = 4
x+ 5 = 9
3x = 12.
The sides of the equilateral triangle are 12 each.